A certain drug can be used to reduce the acid produced by the body and heal damage to the esophagus due to acid reflux. The manufacturer of the drug claims that more than 93% of patients taking the drug are healed within 8 weeks. In clinical​ trials, 215 of 230 patients suffering from acid reflux disease were healed after 8 weeks. Test the​ manufacturer's claim at the α=0.1 level of significance.

A. Find the test statistic, z0

B. Find the P value

C. Choose the correct conclusion below

Do not reject null hypothesis. There is sufficient evidence to conclude more than 93% of patients are taking the drug and healed within 8 weeks

reject null hypothesis. There is sufficient evidence to conclude more than 93% of patients are taking the drug and healed within 8 weeks

reject null hypothesis. There is insufficient evidence to conclude more than 93% of patients are taking the drug and healed within 8 weeks

Do not reject null hypothesis. There is insufficient evidence to conclude more than 93% of patients are taking the drug and healed within 8 weeks
Answer #1

null hypothesis:Ho: p<=0.93

alternate hypothesis: p>0.93

for std error =(p(1-p)/n)1/2 =0.0169

phat=216/229 =0.9432

test stat Z=(phat-p)/std error =0.7848

crtitical value =1.645

C.Do not reject the null hypothesis. There is insufficient evidence to conclude that more than 93​% of patients taking the drug are healed within 8 weeks.